OpenStudy (anonymous):

Which of these is a simplified form of the equation 9p + 3 = − p + 7 + 2p + 3p? 5p = 4 15p = 10 9 = 4 12 = 13

OpenStudy (anonymous):

use mathway.com for an easier way of doing this :)

OpenStudy (anonymous):

you have to pay for that.

OpenStudy (anonymous):

no you dont.

OpenStudy (anonymous):

okay take your equation. 9p+3=-p+7+2p+3p what do you do now.?

OpenStudy (anonymous):

you do. i just asked it a question, and it told me to make an account. and i did. but then it asked me to enter my credit card number.

OpenStudy (anonymous):

hm, never does that to me. i use it all the time.

OpenStudy (anonymous):

how do you use it?

OpenStudy (anonymous):

would you like me to show you how to solve this.?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

take your equation 9p+3=-p+7+2p+3p now tell me what you think you need to do now

OpenStudy (anonymous):

9p + 3? You would have to add that first right?

OpenStudy (anonymous):

no no no. your simplifying. think "simple"

OpenStudy (anonymous):

let me give you a hint

OpenStudy (anonymous):

okay, thank you..

OpenStudy (anonymous):

9p+3=-p+7+5p what looks different here than from the original question.?

OpenStudy (anonymous):

different numbers..

OpenStudy (anonymous):

but simplified.

OpenStudy (anonymous):

true, but how did the numbers change.?

OpenStudy (anonymous):

they were added?

OpenStudy (anonymous):

there you go.

OpenStudy (anonymous):

so would it be a?

OpenStudy (anonymous):

when you simplify the right side more it will look like this 9p+3=4p+7 now the numbers need to be moved. So you subtract 4p from both sides to get 5p+3=7 Then you subtract 3 from both sides to get 5p=4 so yes, A (:

OpenStudy (anonymous):

oh goodness, thank you so much! :)

OpenStudy (anonymous):

i really appreciate you explaining it to me :)

OpenStudy (anonymous):

no problem. Remember, think "simple"

OpenStudy (anonymous):

i will :)

OpenStudy (anonymous):

have a good night (:

OpenStudy (anonymous):

you as well. :)